find the domain and range of each function. write your answer in interval notation. F(x)= x^2-5
Answer by lwsshak3(11628)

find the domain and range of each function. write your answer in interval notation. F(x)= x^2-5
***
Domain; all real numbers, (-∞,∞)
Range:
express x in terms of y
y=x^2-5
x^2=y+5
x=±√(y+5)
y+5≥0
y≥-5
Range: [-5,∞)